OpenStudy (anonymous):

Can you help me solve this system: 15k^2 + 24k + 8kn + n^2 - 6n + 8 = 0 8k^2 - 3k - 4kn - n^2 - 2n + 12 = 0

OpenStudy (anonymous):

\[15k^{2} + 24k + 8kn + n^{2} - 6n + 8 = 0\] \[8k^{2} - 3k - 4kn - n^{2} - 2n + 12 = 0\]

OpenStudy (anonymous):

Ok, but how do you solve it algebraically?

OpenStudy (anonymous):

Are you sure that your equations are correct? I do not believe there is an easy way to do it by hand

OpenStudy (anonymous):

Well, even if the coefficients aren't correct (though I believe they are), do you know the algorithm for solving these kinds of equations?

OpenStudy (anonymous):

Or could you help me with the original problem: Finding the common tangent lines to circles \[(x-4)^{2} + (y - 3)^{2} = 1\] \[(x-2)^{2} + (y + 1)^{2} = 9\]

OpenStudy (anonymous):

I substituted y = kx + n to both of theses equations and got that system with k and n. Is there some other way to find those tangent lines?

OpenStudy (anonymous):

*these

OpenStudy (anonymous):

That's great thank you :)

OpenStudy (anonymous):

yw